TATP (triacetone triperoxide)

1) Assuming the small stable product molecules are gaseous water, carbon dioxide, and nitrogen, write the balanced chemical equation that illustrates your explosives decomposition. On a clean sheet of paper, show the balanced chemical equation using the lewis structures for each reactant and product (one Lewis structure of each product with its stoichiometric coefficient is sufficient), include lone pairs of electrons where appropriate. Using hand drawn figures is acceptable. Use whole numbers for your stoichiometric coefficients, adding in oxygen as a reactant or product, as necessary, in the form of O2(g).

2) Calculate the volume of gas that your explosive would produce if 1 g of it was used. Assume standard temperature and pressure (STP) and any water produced is in the gaseous state.
Answer #1

1)

The structure of TATP (triacetone triperoxide) is

Since it contains only C, H and O, the complete decomposition will result in only carbon dioxide and water.

Using the chemical formula, we can write the following equation for its decomposition.

C_9H_{18}O_6 + O_2 \rightarrow CO_2 + H_2O

Note that all the carbon and hydrogens in the product side must come from TATP itself.

Hence, to balance the reaction

* To balance the number of carbon on both side we multiply 9 on CO2 and to balance the number of hydrogens on both side we multiply 9 on H2O

  C_9H_{18}O_6 + O_2 \rightarrow 9CO_2 + 9H_2O

* Now there are 9 C and 18 H on both side of the equation. There are 8 O on reactant side and 27 O on product side.

Hence, we multiply 21 on O2 to balance the reaction.

21

Now, to make all the stoichiometric coefficients integers, we multiply 2 throughout.

Hence, the balanced reaction will be

2C_9H_{18}O_6 + 21O_2_{(g)} \rightarrow 18CO_2_{(g)} +18H_2O_{(g)}

As Lewis structures we can write the balanced reaction as follows

2)

Molecular mass of TATP can be calculated as

9 × C+18 × H 6 × O = 9 × 12 g/mol+18 × 1 g/mol + 6 × 16 g/mol 108 g/mol 18 g/mol 96 g/mol 222 g/mol

Hence, the number of moles of TATP in 1 g is

mass ะ 0.0045 mol molarmass 222 g/mol

Now, according to the balanced reaction, 2 moles of TATP produces 18 moles of CO2 and 18 moles of H2O gases.

Hence, 2 moles of TATP produces a total of 18+18=36 moles of gas.

Hence, 0.0045 moles of TATP will produce

× 0.0045 mol TATP = 0.081 mol 2 mol TATP

Hence, the amount of gases formed from the explosion is 0.081 mol.

At STP, T = 273.15 K and P = 1 bar

Hence, the volume of gases produced can be calculated using the ideal gas equation

PV = nRT

Hence, the volume of 0.081 mol of gas produced can be calculated as

0.081 mol × 8.314 × 10-2 Lbar K-1 mol-1 × 273.15 K nRT 1 bar IT + VA 1.84 L

Hence, the volume of gas produced from the explosion of 1 g of TATP is about 1.84 L.